OpenStudy (anonymous):

log (base 8) 2 + log (base 4) 2

OpenStudy (candy13106):

hey keith can u plz help me on my question

OpenStudy (anonymous):

So we have \[log_8(2)+log_4(2)\] We can use the property \(log_a(b)=\frac{log_{10}(b)}{log_{10}(a)}\) And simplify the equation to: \[log_8(2)+log_4(2)=\frac{log_{10}(2)}{log_{10}(8)}+\frac{log_{10}(2)}{log_{10}(4)}\]

OpenStudy (candy13106):

hey keith can u plz help me on my question

OpenStudy (candy13106):

im kandy44

OpenStudy (candy13106):

but it got off conection so now im on my old account

OpenStudy (candy13106):

so plz help me keith

OpenStudy (anonymous):

Thanks man! Really appreciate it.

OpenStudy (candy13106):

keith

OpenStudy (anonymous):

Alright Candy. tag me in the question by typing @keithafascalclover

OpenStudy (anonymous):

And @dckovach you can simplify the question a little more haha but ill let ya do that. You good?

OpenStudy (candy13106):

ok its on kandy44

OpenStudy (candy13106):

ok

OpenStudy (candy13106):

u got to find it somwe where at the bottom

OpenStudy (anonymous):

You cant tag me?

OpenStudy (candy13106):

wat does tag mean

OpenStudy (anonymous):

Yeah, I can figure it out from here! Thanks for taking the time to solve that equation.

OpenStudy (anonymous):

question, not equation